The sum of the measure of the interior angles of a polygon is 1980°.
What polygon is it?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

The sum of the measure of the interior angles of a polygon is 1980°.
-----
(n-1)*180 = 1980
-----
n-1 = 11
----
n = 12
-----
Ans: duodecagon
=============
Cheers,
Stan H.
=============

Answer by Alan3354(69443)

The sum of the measure of the interior angles of a polygon is 1980°.
What polygon is it?
--------------
1980 = 180*(n-2)
n-2 = 11
n = 13 sides
--> 13agon, there might be a name for it.